OpenStudy (anonymous):

4x-y=2 9x-2y=7 Use the method of substitution to solve the system of linear equations

OpenStudy (tkhunny):

Well, do it! Solve the top one for 'y', since that looks easiest, and substitute away!

OpenStudy (anonymous):

i have been at this assignment since 7 this morning

OpenStudy (tkhunny):

Take a break. Why would you do that? So, what's stopping you from doing this one?

OpenStudy (anonymous):

all the numbers are looking the same, The assignment is due at 2 am my time

OpenStudy (tkhunny):

Okay, solve for y: 4x-y=2

OpenStudy (anonymous):

1y=-2+4x

OpenStudy (anonymous):

y=-2+4x

OpenStudy (anonymous):

9x=-2(-2+4x)=7

OpenStudy (tkhunny):

It helps to avoid errors by using various algebraic properties. -2 + 4x is okay, but that leading negative sign is often overlooked. I would have been happier with: y = 4x - 2 9x-2(4x-2)=7

OpenStudy (anonymous):

9x-8x+4=7

OpenStudy (tkhunny):

Almost there!

OpenStudy (anonymous):

1x+4=7

OpenStudy (anonymous):

1x+4+4=7+4 1x=11

OpenStudy (tkhunny):

No good. "-4", not "+4"

OpenStudy (anonymous):

1x+4-4=7-4 1x=3

OpenStudy (tkhunny):

No point keeping that '1' around on the front. x = 3. Now, go use an original equation and find the value for 'y'.

OpenStudy (anonymous):

4(3)-y=2 12-y=2

OpenStudy (anonymous):

y=2-12 y=6

OpenStudy (tkhunny):

?? 2 - 12 = ??

OpenStudy (anonymous):

y=10

OpenStudy (tkhunny):

No, that's 12 - 2.

OpenStudy (anonymous):

-10

OpenStudy (tkhunny):

x = 3 y = -10 Does that work in the second equation?